# SemHARQ: Semantic-Aware HARQ for Multi-task Semantic Communications

### v1

@@ III-A Feature distortion evaluation (FDE) network
For the backward propagation, simply back-propagating through the binarization function as if it had been an identity function may cause discrepancies between the forward and backward pass, leading to high variance. Therefore, to improve the efficiency and stability of gradient estimation, Gumbel-softmax distributions are used, where each sample is a differentiable proxy of the corresponding discrete sample. Furthermore, note that straightforwardly applying single-dimension Gumbel-softmax is probable to sample repeatedly since the $R$ selections act independently and are not aware of each other. To solve this issue, a multi-dimension Gumbel-softmax scheme is proposed, where $R$ Gumbel-softmax distributions are designed simultaneously from $R$ categorical distributions generating from ${\bm{{d}}}_{j}$. The $R$ Gumbel-softmax distributions are used to estimate the gradients of $R$ samplings in the forward propagation.
@@
Specifically, a categorical vector ${\bm{{v}}}_{j}^{r}\in\mathbb{R}^{B}$ is designed for the $r$-th sampling, where $r=1,...,R$. It is obtained from ${\bm{{d}}}_{j}^{r}\in\mathbb{R}^{B}$ after a softmax function and can be denoted as
${\bm{{v}}}_{j}^{r}=\mathrm{softmax}({\bm{{d}}}_{j}^{r}),$ (10)
@@
where ${\bm{{d}}}_{j}^{r}$ is obtained after replacing the largest $r-1$ values with 00 in ${\bm{{d}}}_{j}$. On this basis, ${\bm{{v}}}_{j}^{r}$ serves as a probability vector involving the variables in ${\bm{{d}}}_{j}^{r}$, and can be utilized to form the $r$-th Gumbel-softmax distribution. Since the samples from Gumbel-softmax distributions will become identical to the samples from the corresponding categorical distributions along with training, the maximum value in each categorical vector will most likely be sampled, matching the forward propagation. Therefore, ${\bm{{v}}}_{j}^{r}$ is used to select the $r$-th largest value in ${\bm{{d}}}_{j}$, mitigating repetitive samplings. Consequently, the sample vector $\bm{w}_{j}^{r}$ in $r$-th sampling is represented as
$\begin{split}{\bm{{w}}}_{j}^{r}=\mathrm{softmax}\biggl{(}\Bigl{(}\log({\bm{{v}}}_{j}^{r})+\bm{g}^{r}\Bigl{)}/\tau\biggl{)},\end{split}$ (11)
@@
where $\bm{g}^{r}$ is a sampling vector drawn from Gumbel(0,1) distribution. The softmax function in (11) is used as a continuous, differentiable approximation to $\arg\max$, and $\tau$ serves as the softmax temperature. As $\tau$ approaches 00, the distribution will become more discrete and the sample vectors will converge to one-hot vectors. During training, the temperature parameter is decreased by an exponentially decreasing curve as in [34, 35]. Finally, all the $R$ sample vectors are summed up to be a continuous relaxation of the quantized ${\bm{{p}}}_{j+1}$ for backward propagation, which is denoted as
$\bm{w}_{j}=\sum_{r=1}^{R}{\bm{w}}_{j}^{r}.$ (12)
@@ III-B Feature importance ranking (FIR)
To identify the importance levels of different features in $\bm{f}$ w.r.t. specific tasks, we refer to gradient-weighted class activation mapping (Grad-CAM) [36, 37], where visual explanations of the decision-making process of a convolutional neural network (CNN) are proposed. In Grad-CAM, the important regions within an input image are identified by leveraging the gradients flowing into the last convolutional layer of the CNN. The gradients represent the sensitivity of each feature map w.r.t. the target class. The higher the gradient value, the more influential the corresponding feature map is for the prediction.
@@
Figure 5: The framework of feature importance ranking (FIR), where features with deeper colors in $\bm{f}$ have higher importance levels.
@@ IV Semantic-aware HARQ
In this section, the process of SemHARQ is further analyzed. Meanwhile, the training strategy of SemHARQ is proposed to optimize the training process.
@@ IV-A SemHARQ
$\bm{\widetilde{z}}_{0}={f}_{\text{sfs}}\biggl{(}\emph{C}_{\alpha}\Bigl{(}\emph{S}_{\beta}(\bm{x})\Bigl{)}\biggl{)},$ (15)
@@
where $\emph{S}_{\beta}(\cdot)$ and $\emph{C}_{\alpha}(\cdot)$ denote the multi-task semantic encoder with the parameter set $\beta$ and the JSC-Encoder with the parameter set $\alpha$, respectively, and $f_{\text{sfs}}(\cdot)$ indicates the scalable feature selector, whose structure is shown in Fig. 7. Consequently, $\bm{\widetilde{z}}_{0}$ consists of $B$ features chosen from the encoded feature vector according to the importance measure vector Ind and the given channel condition. After normalizing and passing through the channel, the received feature vector is represented as $\hat{\bm{z}}_{0}$ at the receiver.
Figure 6: The structure of SemHARQ.
@@
Then, $\hat{\bm{z}}_{0}$ is detected by the retransmission identification module to determine the need for retransmission. Specifically, a retransmission criterion based on the received data uncertainty is applied [12]. Consider the acquisition of a concatenated feature vector $\overline{\bm{z}}_{j}$ after $j$ retransmissions, the receiver repeatedly requests the transmitter to retransmit until
$\operatorname{\overline{SNR}}>\min\biggl{(}\theta_{0}\mathcal{L}\Bigl{(}\mathcal{U}({\bm{\overline{z}}_{j}})\Bigl{)},\theta_{\mathrm{SNR}}\biggl{)},$ (16)
@@
where $\operatorname{\overline{SNR}}$ stands for the receive-SNR after the previous transmissions, $\mathcal{U}(\cdot)$ denotes the received information uncertainty, $\mathcal{L}(\cdot)$ represents a monotonically increasing reshaping function which is set to $\mathcal{L}(x)=1+x$, $\theta_{0}$ indicates a given conversion ratio between the uncertainty and SNR, and $\theta_{\mathrm{SNR}}$ is a given maximum SNR. Particularly, the receive-SNR is obtained by integrating the information after multiple retransmissions, which is given by
$\operatorname{\overline{SNR}}=\frac{2P}{\sigma^{2}}\sum_{i=1}^{J}\left|h^{(i)}\right|^{2},$ (17)
@@
where $P$ is the average power of the received semantics, $h^{(i)}$ represents the channel coefficient at $i$-th retransmission, $\frac{\sigma^{2}}{2}$ indicates the noise variance, $J$ denotes the total number of transmissions. As for the uncertainty, $\mathcal{U}({\bm{\overline{z}}_{j}})$ is measured by the average entropy of the posterior probability distribution corresponding to different tasks. In terms of the $k$-th task, the uncertainty is given by
$\mathcal{U}_{k}({\bm{\overline{z}}_{j}})=-\sum_{y_{k}}p_{\gamma}(y_{k}\mid{\bm{\overline{z}}_{j}})\log p_{\gamma}(y_{k}\mid{\bm{\overline{z}}_{j}}),$ (18)
@@
where $y_{k}$ denotes a class label and $\gamma$ is the set of parameters of the decoders and the task performer to be learned. The uncertainty measures the contribution of the received semantics to the task performance from the perspective of information entropy, and thus serves as a criterion for measuring the adequacy of the amount of received information. Furthermore, since the uncertainty will be constantly bridged by increasing the receive-SNR or equivalently the number of retransmissions, the inequality in (16) is intuitively used to evaluate the adequacy of the number of retransmissions.
@@
After passing through the retransmission identification module, if $\hat{\bm{z}}_{0}$ passes the detection successfully, it is fed into the decoders and the multiple task performers to generate the multi-task predictions. The received feature vector is decoded as
${\bm{c}_{k}}=\emph{T}_{\phi_{k}}\biggl{(}\emph{S}^{-1}_{\beta_{k}^{-1}}\Bigl{(}\emph{C}^{-1}_{\alpha^{-1}}({\hat{\bm{z}}_{0}})\Bigl{)}\biggl{)},$ (19)
@@
where $\bm{c}_{k}$ denotes the predicted probability vector of the $k$-th task, $\emph{C}^{-1}_{\alpha^{-1}}(\cdot)$, $\emph{S}^{-1}_{\beta_{k}^{-1}}(\cdot)$ and $\emph{T}_{\phi_{k}}(\cdot)$ represent the JSC-Decoder with parameter set ${\alpha^{-1}}$ , the semantic decoder of the $k$-th task with parameter set ${\beta_{k}^{-1}}$ and the task performer of the $k$-th task with parameter set ${\phi_{k}}$, respectively. Additionally, if $\hat{\bm{z}}_{0}$ fails to pass the detection, the FDE network is utilized to generate the feedback vector $\bm{p}_{1}$ to trigger the first distortion-aware retransmission. The distorted features in $\hat{\bm{z}}_{0}$ are required to be retransmitted according to the distortion evaluation. In particular, the incremental transmission of new features in order of importance levels is combined with the retransmission, further improving the transmission efficiency. Furthermore, after the first retransmission, $\hat{\bm{z}}_{1}$ is obtained at the receiver and concatenated with the cached semantics from the previous transmission. The retransmission process will continue until either the receive-SNR is sufficiently large or the predetermined maximum number of retransmissions is reached.
Figure 7: The structure of the scalable feature selector for HARQ, where the semantic features are transmitted directly after scalable feature selection at the first transmission.
@@
2:$\xi_{\mathcal{F}}$.
3:Sample from $(\bm{s},\bm{\hat{s}})$, $\bm{s}$ and $\bm{\hat{s}}$, and feed the samples into $\mathcal{F}$.
4:Estimate the mutual information of $\bm{s}$ and $\bm{\hat{s}}$ by (7).
5:Train $\xi_{\mathcal{F}}$ by loss function $-I(\bm{s};\bm{\hat{s}}|\mathcal{F})$.
@@
1:Images $\bm{x}$ from training dataset, initialized parameters $\alpha$, $\beta$, $\alpha^{-1}$, $\beta_{k}^{-1}$, $\phi_{k}$ with $\bm{f}$ fully transmitted; $k=1,...,K$.
2:$\emph{S}_{\beta}(\cdot)$, $\emph{C}_{\alpha}(\cdot)$, $\emph{C}_{\alpha^{-1}}^{-1}(\cdot)$, $\emph{S}^{-1}_{\beta_{k}^{-1}}(\cdot)$, $\emph{T}_{\phi_{k}}(\cdot)$.
@@
5:     Obtain $\bm{p}_{j}$ from the FDE network and
6:     Send $\bm{p}_{j}$ back to the transmitter.
7:     Trigger distortion-aware retransmission according to $\bm{p}_{j}$.
@@
9:Receiver: ${\bm{\overline{z}}}_{j}=[{\bm{\hat{z}}}_{1},...,{\bm{\hat{z}}}_{j}]$, ${\bm{c}_{k}}=\emph{T}_{\phi_{k}}\Bigl{(}\emph{S}^{-1}_{\beta_{k}^{-1}}(\emph{C}^{-1}_{\alpha^{-1}}(\bm{\overline{z}}_{j}))\Bigl{)}$.
10:Train $\alpha$, $\beta$, $\alpha^{-1}$, $\beta_{k}^{-1}$, $\phi_{k}$ by loss function $\emph{L}_{\text{E2E}}$ ; $k=1,...,K$.
@@ IV-B Training strategy
To improve the efficiency of SemHARQ training, a three-stage training strategy is proposed. We focus on three representative tasks in intelligent traffic scenarios for experiments, namely vehicle ReID, vehicle color classification, and vehicle type classification. DenseNet-121 pre-trained on ImageNet is used as the initial model. In the first stage, to improve the performance of feature distortion evaluation, the FDE network is trained with other trainable modules frozen. The training loss can be expressed as,
@@
$L_{\text{MI}}=-I(\bm{s};\bm{\hat{s}}|\mathcal{F}),$ (22)
where $I(\bm{s};\bm{\hat{s}}|\mathcal{F})$ is calculated in (7).
Then, to calculate the feature importance levels, the whole system is trained with the entire $\bm{f}$ transmitted without considering the constraint of communication capacity. The overall E2E loss is represented by
@@
where $L_{r}$, $L_{c}$, and $L_{t}$ correspond to the ReID, color classification, and type classification losses, respectively, $\lambda_{r}$, $\lambda_{c}$, and $\lambda_{t}$ denote the respective task weights. Specifically, $L_{r}$ consists of the hard-mining triplet loss [38] and the cross-entropy loss, defined as $L_{r}=\mathcal{L}_{ht}(a,p,n)+\mathcal{L}_{ce}(\bm{y},\bm{\hat{y}})$, where $a,p$ and $n$ represent anchor, positive and negative elements, respectively. The hard-mining triplet loss $\mathcal{L}_{ht}(a,p,n)$ is defined as
@@
where $d_{ap}$ and $d_{an}$ denote the distances between the extracted features of the anchor and positive/negative images. $\mathcal{L}_{ce}(\bm{y},\bm{\hat{y}})$ indicates the cross-entropy loss, which is defined as
@@
Output
($j$-th retransmission)
$\bm{\widetilde{z}}_{j}$ $B$ /
@@
FDE network ($j$-th retransmission) $\mathcal{F}$
@@
Finally, the SemHARQ considering scalable feature selection is trained based on Ind with limited bandwidth and time slot, where all the trainable parameters are optimized. The most important $B$ features are transmitted, where $B$ is determined by channel conditions. The pseudocode of the whole training process is shown in Algorithm 1.

### v2

@@ III-A FDE network
For the backward propagation, simply back-propagating through the binarization function as if it had been an identity function may cause discrepancies between the forward and backward pass, leading to high variance. Therefore, to improve the efficiency and stability of gradient estimation, Gumbel-softmax distributions are used, where each sample is a differentiable proxy of the corresponding discrete sample. Furthermore, note that straightforwardly applying single-dimension Gumbel-softmax is probable to sample repeatedly since the $R$ selections act independently and are not aware of each other. To solve this issue, a multi-dimension Gumbel-softmax scheme is proposed, where $R$ Gumbel-softmax distributions are designed simultaneously from $R$ categorical distributions generating from ${\bm{{d}}}^{j}$. The $R$ Gumbel-softmax distributions are used to estimate the gradients of $R$ samplings in the forward propagation.
@@
Specifically, a categorical vector ${\bm{{v}}}^{jr}\in\mathbb{R}^{B}$ is designed for the $r$-th sampling, where r=1,,Rr=1,...,R. It is obtained from ${\bm{{d}}}^{jr}\in\mathbb{R}^{B}$ after a softmax function and can be denoted as
${\bm{{v}}}^{jr}=\mathrm{softmax}({\bm{{d}}}^{jr}),$ (13)
@@
where ${\bm{{d}}}^{jr}$ is obtained after replacing the largest r1r-1 values with 0 in ${\bm{{d}}}^{j}$. On this basis, ${\bm{{v}}}^{jr}$ serves as a probability vector involving the variables in ${\bm{{d}}}^{jr}$, and can be utilized to form the $r$-th Gumbel-softmax distribution. Since the samples from Gumbel-softmax distributions will become identical to the samples from the corresponding categorical distributions along with training, the maximum value in each categorical vector will most likely be sampled, matching the forward propagation. Therefore, ${\bm{{v}}}^{jr}$ is used to select the $r$-th largest value in ${\bm{{d}}}^{j}$, mitigating repetitive samplings. Consequently, the sample vector $\bm{w}^{jr}$ in $r$-th sampling is represented as
$\begin{split}{\bm{{w}}}^{jr}=\mathrm{softmax}\biggl{(}\Bigl{(}\log({\bm{{v}}}^{jr})+\bm{g}^{r}\Bigl{)}/\tau\biggl{)},\end{split}$ (14)
@@
where $\bm{g}^{r}$ is a sampling vector drawn from Gumbel(0,1) distribution. The softmax function in Eq. (14) is used as a continuous, differentiable approximation to $\arg\max$, and $\tau$ serves as the softmax temperature. As $\tau$ approaches 0, the distribution will become more discrete and the sample vectors will converge to one-hot vectors. During training, the temperature parameter is decreased by an exponentially decreasing curve as in [45, 46]. Finally, all the $R$ sample vectors are summed up to be a continuous relaxation of the quantized ${\bm{{p}}}^{(j+1)}$} for backward propagation, which is denoted as
$\bm{w}^{j}=\sum_{r=1}^{R}{\bm{w}}^{jr}.$ (15)
@@ III-B Feature importance ranking (FIR)
To identify the importance levels of different features in $\bm{f}$ w.r.t. specific tasks, we refer to gradient-weighted class activation mapping (Grad-CAM) [47], where visual explanations of the decision-making process of a CNN are proposed. In Grad-CAM, the important regions within an input image are identified by leveraging the gradients flowing into the last convolutional layer of the CNN. The gradients represent the sensitivity of each feature map w.r.t. the target class. The higher the gradient value, the more influential the corresponding feature map is for the prediction.
@@
Figure 5: The framework of FIR, where features with deeper colors in $\bm{f}$ attach higher importance.
@@ IV Semantic-aware HARQ
In this section, the process of SemHARQ is further analyzed. Note that to transmit the meanings of the source efficiently, all the functions of SemHARQ are performed at the level of semantic features. The JSCC is utilized in our framework to encode semantic features directly into channel transmission symbols. Furthermore, the training strategy of SemHARQ is proposed to optimize the training process.
@@ IV-A SemHARQ
$\bm{\widetilde{z}}^{0}={f}_{\text{sfs}}\biggl{(}\emph{C}_{\alpha}\Bigl{(}\emph{S}_{\beta}(\bm{x})\Bigl{)}\biggl{)},$ (18)
@@
where $\emph{S}_{\beta}(\cdot)$ and $\emph{C}_{\alpha}(\cdot)$ denote the multi-task semantic encoder with the parameter set $\beta$ and the JSC-Encoder with the parameter set $\alpha$, respectively, and $f_{\text{sfs}}(\cdot)$ indicates the scalable feature selector, whose structure is shown in Fig. 7. Consequently, $\bm{\widetilde{z}}^{0}$ consists of $B$ features chosen from the encoded feature vector according to the importance measure vector Ind and the given channel condition. After normalizing and passing through the channel, the received feature vector is represented as $\hat{\bm{z}}^{0}$ at the receiver.
Figure 6: The working flow of SemHARQ.
@@
Then, $\hat{\bm{z}}^{0}$ is detected by the retransmission identification module to determine the need for retransmission. Specifically, a retransmission criterion based on the received data uncertainty is applied [19]. Consider the acquisition of a concatenated feature vector $\overline{\bm{z}}^{j}$ after $j$ retransmissions, the receiver repeatedly requests the transmitter to retransmit until
$\operatorname{\overline{SNR}}>\min\biggl{(}\theta_{0}\mathcal{L}\Bigl{(}\mathcal{U}({\bm{\overline{z}}^{j}})\Bigl{)},\theta_{\mathrm{SNR}}\biggl{)},$ (19)
@@
where $\operatorname{\overline{SNR}}$ stands for the receive-SNR after the previous transmissions, $\mathcal{U}(\cdot)$ denotes the received information uncertainty, $\mathcal{L}(\cdot)$ represents a monotonically increasing reshaping function which is set to $\mathcal{L}(x)=1+x$, $\theta_{0}$ indicates a given conversion ratio between the uncertainty and SNR of the combined information, and $\theta_{\mathrm{SNR}}$ is a given maximum receive-SNR. Particularly, the receive-SNR is obtained by integrating the information after multiple retransmissions, which is given by
$\operatorname{\overline{SNR}}=\frac{2P}{\sigma^{2}}\sum_{j=0}^{J}\left\|\bm{h}^{j}\right\|^{2},$ (20)
@@
where $P$ is the average power of the semantics, $\bm{h}^{j}$ represents the channel coefficient vector at the $j^{th}$ retransmission, $\frac{\sigma^{2}}{2}$ indicates the noise variance, $J$ denotes the total number of retransmissions. As for the uncertainty, $\mathcal{U}({\bm{\overline{z}}^{j}})$ is measured by the average entropy of the posterior probability distribution corresponding to different tasks. In terms of the $k$-th task, the uncertainty is given by
$\mathcal{U}_{k}({\bm{\overline{z}}^{j}})=-\sum_{y_{k}}p_{\gamma}(y_{k}\mid{\bm{\overline{z}}^{j}})\log p_{\gamma}(y_{k}\mid{\bm{\overline{z}}^{j}}),$ (21)
@@
where $y_{k}$ denotes a class label and $\gamma$ is the set of parameters of the decoders and the task performer to be learned. The uncertainty measures the contribution of the received semantics to the task performance from the perspective of information entropy, and thus serves as a criterion for measuring the adequacy of the amount of received information. Furthermore, since the uncertainty will be constantly bridged by increasing the receive-SNR or equivalently the number of retransmissions, the inequality in Eq. (19) is intuitively used to evaluate the adequacy of the number of retransmissions.
@@
After passing through the retransmission identification module, if $\hat{\bm{z}}^{0}$ passes the detection successfully, it is fed into the decoders and the multiple task performers to generate the multi-task predictions. The received feature vector is decoded as
${\bm{c}_{k}}=\emph{T}_{\phi_{k}}\biggl{(}\emph{S}^{-1}_{\beta_{k}^{-1}}\Bigl{(}\emph{C}^{-1}_{\alpha^{-1}}({\hat{\bm{z}}^{0}})\Bigl{)}\biggl{)},$ (22)
@@
where $\bm{c}_{k}$ denotes the predicted probability vector of the $k$-th task, $\emph{C}^{-1}_{\alpha^{-1}}(\cdot)$, $\emph{S}^{-1}_{\beta_{k}^{-1}}(\cdot)$ and $\emph{T}_{\phi_{k}}(\cdot)$ represent the JSC-Decoder with parameter set $\alpha^{-1}$} , the semantic decoder of the $k$-th task with parameter set $\beta_{k}^{-1}$} and the task performer of the $k$-th task with parameter set $\phi_{k}$}, respectively. Additionally, if $\hat{\bm{z}}^{0}$ fails to pass the detection, the FDE network is utilized to generate the feedback vector $\bm{p}^{1}$ to trigger the first distortion-aware retransmission. The distorted features in $\hat{\bm{z}}^{0}$ are required to be retransmitted according to the distortion evaluation. Note that the number of transmitted features at every transmission is adaptive to the channel conditions. Therefore, the incremental transmission of new features in descending order of importance levels is combined with the retransmission if there are remaining channel resources. On this basis, the transmission robustness and efficiency are enhanced simultaneously. Furthermore, after the first retransmission, $\hat{\bm{z}}^{1}$ is obtained at the receiver and concatenated with the cached semantics from the previous transmission. The retransmission process will continue until either the receive-SNR is sufficiently large or the predetermined maximum number of retransmissions is reached.
Figure 7: The structure of the scalable feature selector for SemHARQ.
@@
[…]
4:Estimate the mutual information of $\bm{s}$ and $\bm{\hat{s}}$ by Eq. (7).
[…]
@@
1:Images $\bm{x}$ from training dataset, initialized parameters $\alpha$, $\beta$, $\alpha^{-1}$, $\beta_{k}^{-1}$, $\phi_{k}$ with $\bm{f}$ fully transmitted; k=1,,Kk=1,...,K.
[…]
@@
5:  Obtain $\bm{p}^{j}$ from the FDE network and
6:  Send $\bm{p}^{j}$ back to the transmitter.
7:  Trigger distortion-aware retransmission according to $\bm{p}^{j}$.
@@
9:Receiver: ${\bm{\overline{z}}}^{j}=[{\bm{\hat{z}}}^{1},...,{\bm{\hat{z}}}^{j}]$, ${\bm{c}_{k}}=\emph{T}_{\phi_{k}}\Bigl{(}\emph{S}^{-1}_{\beta_{k}^{-1}}(\emph{C}^{-1}_{\alpha^{-1}}(\bm{\overline{z}}^{j}))\Bigl{)}$.
10:Train $\alpha$, $\beta$, $\alpha^{-1}$, $\beta_{k}^{-1}$, $\phi_{k}$ by loss function $\emph{L}_{\text{E2E}}$ ; k=1,,Kk=1,...,K.
@@ IV-B Training strategy
To improve the efficiency of SemHARQ training, a three-stage training strategy is proposed. We focus on three representative tasks in intelligent traffic scenarios for experiments, namely vehicle ReID, vehicle color classification, and vehicle type classification. DenseNet-121 pre-trained on ImageNet is used as the initial model [48]. In the first stage, to improve the performance of feature distortion evaluation, the FDE network is trained with other trainable modules frozen. The training loss can be expressed as,
@@
[…]
Then, in the second stage, to calculate the feature importance levels, the whole system is trained with the entire $\bm{f}$ transmitted without considering the constraint of communication capacity. The overall E2E loss is represented by
@@
[…]
@@
[…]
@@
Output ($j^{th}$ retransmission)
$\bm{\widetilde{z}}^{j}$ $B$ /
@@
FDE network
($j^{th}$ retransmission)
$\mathcal{F}$
@@
Finally, the SemHARQ considering scalable feature selection is trained based on the importance measure vector Ind with limited bandwidth and time slot, where all the trainable parameters are optimized. The most important $B$ features are transmitted, where $B$ is determined by channel conditions. The pseudocode of the whole training process is shown in Algorithm 1.